# Dressing rotating black holes with anisotropic matter

### v2

@@ Abstract
We present a new rotating black hole solution to the Einstein equations as an extension of the Kerr spacetime. To derive this solution, we use the Newman-Janis algorithm and results from the non-complexification procedure. Interestingly, the solution we find may not be uniquely characterized by asymptotic parameters such as mass, angular momentum, and charge, thereby challenging the no-hair theorem. We also analyze in detail how this additional characteristics (“hair”) affects the thermodynamic properties of the black hole.
@@ 1 Introduction
A key solution describing a rotating black hole in vacuum is the Kerr metric [8], serves as a foundational reference for deriving more general rotating black hole solutions that incorporate matter fields. The Kerr solution provides a remarkably accurate description of astrophysical black holes and is widely accepted as the best approximation for rotating black holes in the Universe. However, real black holes in galaxies exist in environments filled with dark matter and dark energy. Thus, it is necessary to explore more general solutions that take into account the effects of a cosmological constant and matter fields. Studying such extensions can offer deeper insights into the dynamics and observational signatures of astrophysical black holes.
This naturally raises the question: How can we derive solutions for rotating black holes that coexist with surrounding matter fields? The first major extension in this direction was the Kerr-Newman solution [9], which describes a rotating, charged black hole in Einstein-Maxwell theory . This solution was derived using the Newman-Janis (NJ) algorithm [10], a method that has been widely studied and debated [11, 12, 13, 14, 15, 16, 17]. While the underlying reason how the algorithm works remains unclear, its empirical success makes it a valuable tool for deriving rotating black hole solutions. Recent research has focused on extending this algorithm to incorporate additional matter fields [18, 19, 20, 21]. There is also considerable interest in which types of fields can coexist with rotating black holes while maintaining physical consistency [22, 23, 21]. In various theories of gravity, efforts have been made new rotating black hole solutions in different spacetime dimensions [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38].
@@
The no-hair theorem is well-established for the electrovacuum solution of Einstein-Maxwell theory [82], which states that black holes are characterized solely by their mass, charge, and angular momentum as primary hairs. However, in Einstein-Maxwell theory with additional matter fields beyond ordinary matter or in extensions of general relativity, the applicability of the no-hair theorem is not necessarily evident. When fields other than the electromagnetic field are present, additional hairs may emerge. In the Einstein-scalar theory with minimal coupling, Bekenstein’s no-hair theorem holds under the assumption that the energy density of the scalar field outside the black hole horizon is non-negative [54, 55]. Thus, the theorem can be circumvented if this assumption is violated [56, 57, 58, 59, 60, 62, 63, 64, 65, 66, 67]. Further investigation is required to determine whether such hair should be classified as primary or secondary [68, 69, 70].
@@
In this work, we use results of the “non-complexification procedure” [13], a developments of the Newman-Janis algorithm, as a method for separation of variables for the Einstein equation to explore the structure of axially symmetric, rotating solutions. Let us briefly summarize the strategy to get the rotating solution in this schime. As the base geometry, we consider a static, spherically symmetric metric of the form,
@@
where $d\Omega^{2}=d\theta^{2}+\sin^{2}\theta d\phi^{2}$ represents the angular part of the metric. A black hole solutions in mind, we set $\sigma(r)=r^{2}$. At this stage, we do not presume any specific forms for the functions $f(r)$ and $g(r)$ aside from the asymptotic flatness condition, which requires $f(r)\,,g(r)\to 1$ as $r\to\infty$.
The resulting rotating geometries can be written as the form
@@
In this work, we regard the metric (2) as the starting one to analyze the Einstein equations and get the orthonormal frame. Therefore, we does not assume the predefined static limit (3) but try to get the solution from the general physical requirements. We investigate the structure of these solutions and discuss their implications for black hole physics.
@@ 2 Rotating geometries from the static spherically symmetric black hole
Then, we require the condition that the stress tensor does not have the off-diagonal [^3] Note however that this non-vanishing value does not imply the existence of shear viscosity between the two directions. One could find the orthonormal basis which makes this value vanish by rotating $r$ and $\theta$ directions. $r\theta$ component, $T_{\hat{r}\hat{\theta}}=0$. This is required for fluids undergoing only a rotational motion about a fixed axis (the $z$-axis here), which leads to $R_{\hat{r}\hat{\theta}}=0$ [73]. To calculate the Einstein tensor, we take the orthonormal non-coordinate basis from the metric (6):
@@
$ds^{2}=-\frac{\Delta(r)}{r^{2}}dt^{2}+\frac{r^{2}}{\Delta(r)}dr^{2}+r^{2}d\Omega_{(2)}^{2}\,,$ (13)
where $\Delta=r^{2}-2Mr+v(r)$. The source developing this metric satisfies $T_{\mu\nu}=G_{\mu\nu}/8\pi$ with the Einstein tensor given in Eq. (12). Therefore, we interpret that the metric (10) describes the rotating geometry of this static one.
@@
    where erf$(x)$ denotes the error function. This is an interesting generalization of the Kerr family with a minor modification of the equation of state. Note that $v(r)-rv^{\prime}(r)=v_{2}e^{-v_{c}r^{2}/2}$. Therefore, the stress tensor exponentially decreases with $r$. We analyze the properties of the solution in the next section.
@@
Starting from the metric ansatz (10), we imposes the axially rotating condition $R_{\hat{r}\hat{\theta}}=0$ and select an appropriate equation of state to determine the final form of the metric.
@@ 3 Analysis of the new solution
We now determine the locations of the ergosphere and the event horizon of the rotating black hole. The ergosphere is the region located between the static limit surface and the outer event horizon [78]. The static limit surface is determined at the location $g_{tt}=0$, while the event horizon corresponds to a Killing horizon [79], which is located where $\Delta(r)=0$.
The function $\bar{\Delta}(y)$ monotonically decreases from one at $y=0$ to zero as $y\to\infty$. Expanding $\Delta(r)$ around $r\sim 0$ and $r\gg v_{c},M$, we get
@@
Let us set the radius $r_{c}$ such that $\Delta^{\prime}(r_{c})=0$ At this point, we have
@@
Figure 1: (color online). The characteristic form of the function $\Delta(r)$ for various values of $v_{2}$ [L] and $a$ [R], respectively.
Figure 1 illustrates the shape of $\Delta(r)$ as the function of the coordinate $r$. Figure 1(a) shows the shape of $\Delta(r)$ while varying the value of $v_{2}$, and Fig. 1(b) shows the shape while varying the value of the rotation parameter $a$. We take $M=1$ for simplicity. In Fig. 1(a), the gray curve represents a Schwarzschild black hole with $a=0$ and $v_{2}=0$, while the black curve represents a Kerr black hole with $v_{2}=0$. The red, orange, and green curves represent $\Delta(r)$ with $v_{2}=0.5$, $v_{2}=1.0$, and $v_{2}=1.5$, respectively, with $a=0.3$. In Fig. 1(b), we take $v_{2}=0.5$. The blue dashed line represents $\bar{\Delta}(\sqrt{v_{c}}r)$ with $v_{c}=0.2$. The black curve represents $\Delta(r)$ with $a=0.0$, corresponding to the black hole with anisotropic matter. The red, orange, and green curves represent $\Delta(r)$ with $a=0.3$, $a=0.5$, and $a=0.6$, respectively.
@@
Note that the presence of the $v_{2}\bar{\Delta}(y)$ term always increases the value of $\Delta(r)$. Therefore, the radius of the outer horizon $r_{+}$, given by $\Delta(r)=0$, decreases because of this term. This implies that the $\bar{\Delta}(y)$ term makes the entropy of the black hole decreases making the Kerr black hole has a maximum entropy.
Given the asymptotic values the mass $M$ and the angular momentum parameter $a$, the horizon area varies with $v_{c}$. This makes the thermodynamic law fails [or depends on the non-asymptotic parameter] to be described by the asymptotic values.
In the $a\to 0$ limit, this corresponds to a new static black hole geometry with $\omega_{r}=-1$, and $\omega_{\theta}=\omega_{\phi}=1+v_{c}r^{2}/2$. In other words,
@@ 4 Thermodynamics
We now analyze the shape of the temperature curve for $T_{H}\geq 0$. At large $r_{H}\gg(a^{2}+v_{2}e^{-v_{c}r^{2}_{H}/2})$, the temperature behaves as $T_{H}\propto r^{-1}_{H}$, similar to the Schwarzschild black hole. Conversely, for small $r_{H}$, the term $-(a^{2}+v_{2}e^{-v_{c}r^{2}_{H}/2})$ in the numerator dominates. In this regime, the temperature initially decreases as $T_{H}\propto-(r^{-3}_{H})$ transitioning to $T_{H}\propto-(r^{-1}_{H})$ as $r_{H}$ grows slightly larger. As $r_{H}\to 0$, $T_{H}$ diverges to negative infinity. There exists a specific value of $r_{H}$ where $T_{H}=0$, corresponding to an extremal black hole. Additionally, the $r_{H}$ value at which the black hole temperature reaches its maximum is determined by the condition $dT_{H}/dr_{H}=0$.
@@
$M=\frac{1}{2}\left[(M_{K}+\Phi_{1}C_{1}+\Phi_{2}C_{2})+\sqrt{(M_{K}+\Phi_{1}C_{1}+\Phi_{2}C_{2})^{2}-4\Omega_{H}J(2\Phi_{1}C_{1}+\Phi_{2}C_{2})}\right]\,,$ (23)
where $\Omega_{H}=\frac{a}{r^{2}_{H}+a^{2}}$, $J=Ma$, $S=\pi(r^{2}_{H}+a^{2})$, and $M_{K}=2T_{H}S+2\Omega_{H}J$ denotes the mass of the Kerr black hole.
@@ 5 Summary and discussions
We investigated the Newman-Janis (NJ) algorithm [10] and used results of the “non-complexification procedure” [13] as a method for constructing rotating geometries from static ones. To describe the geometry of a rotating black hole, we solved the Einstein equations, starting with Eq. (2). We first demand the Einstein tensor component $R_{\hat{r}\hat{\theta}}$ vanishes so that the geometry is consistent with the axial symmetry. We then, check the Einstein tensor to get proper equation of states the stress tensor should follow.
However, obtaining a rotating black hole solution becomes more challenging when the NJ algorithm is not well-defined for a given geometry. This issue arises, for instance, when $f(r)\neq g(r)$, when a cosmological constant is present, or when the black hole is immersed in a magnetic field. In the case of $f(r)\neq g(r)$, the standard approach is to solve Einstein’s equations directly by usinging a metric ansatz, and verify whether $R_{\hat{r}\hat{\theta}}=0$. An example of this method applied within the NJ algorithm framework can be found in [85]. For black holes with a cosmological constant, one may follow Carter’s method [86], while black holes immersed in the magnetic field require analysis based on Ernst spacetime [87, 88].

### v3

# CQUeST-2025-0755

@@ Abstract
We present a new rotating black hole solution to the Einstein equations as an extension of the Kerr spacetime. Interestingly, the solution we find may not be uniquely characterized by asymptotic parameters such as mass, angular momentum, and charge, thereby it would be the additional hair. We also analyze in detail how this additional characteristics or this hair affects the thermodynamic properties of the black hole.
@@ 1 Introduction
A key solution describing a rotating black hole in vacuum is the Kerr metric [8], which serves as a foundational reference for deriving more general rotating black hole solutions that incorporate matter fields. The Kerr solution provides a remarkably accurate description of astrophysical black holes and is widely accepted as the best approximation for rotating black holes in the Universe. However, real black holes in galaxies exist in environments filled with dark matter and dark energy. Thus, it is necessary to explore more general solutions that take into account the effects of a cosmological constant and matter fields. Studying such extensions can offer deeper insights into the dynamics and observational signatures of astrophysical black holes.
This naturally raises the question: How can we derive solutions for rotating black holes that coexist with surrounding matter fields? The first major extension in this direction was the Kerr-Newman solution [9], which describes a rotating, charged black hole in Einstein-Maxwell theory. This solution was derived using the Newman-Janis (NJ) algorithm [10], a method that has been widely studied and debated [11, 12, 13, 14, 15, 16, 17]. While the underlying reason how the algorithm works remains unclear, its empirical success makes it a valuable tool for deriving rotating black hole solutions. Recent research has focused on extending this algorithm to incorporate additional matter fields [18, 19, 20, 21]. There is also considerable interest in which types of fields can coexist with rotating black holes while maintaining physical consistency [22, 23, 21]. In various theories of gravity, efforts have been made to introduce new rotating black hole solutions in different spacetime dimensions [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38].
@@
The no-hair theorem is well-established for the electro-vacuum solution of Einstein-Maxwell theory [84], which states that black holes are characterized solely by their mass, charge, and angular momentum as primary hairs. However, in Einstein-Maxwell theory with additional matter fields beyond ordinary matter or in extensions of general relativity, the applicability of the no-hair theorem is not necessarily evident. When fields other than the electromagnetic field are present, additional hairs may emerge. In the Einstein-scalar theory with minimal coupling, Bekenstein’s no-hair theorem holds under the assumption that the energy density of the scalar field outside the black hole horizon is non-negative [54, 55]. Thus, the theorem can be circumvented if this assumption is violated [56, 57, 58, 59, 60, 62, 63, 64, 65, 66, 67]. Further investigation is required to determine whether such hair should be classified as primary or secondary [68, 69, 70].
@@
In this work, we use results of the “non-complexification procedure” [13], a developments of the Newman-Janis algorithm, as a method for separation of variables for the Einstein equation to explore the structure of axially symmetric, rotating solutions. Let us briefly summarize the strategy to get the rotating solution in this scheme. As the base geometry, we consider a static, spherically symmetric metric of the form,
@@
where $d\Omega^{2}=d\theta^{2}+\sin^{2}\theta d\phi^{2}$ represents the angular part of the metric. With black hole solutions in mind, we set $\sigma(r)=r^{2}$. At this stage, we do not presume any specific forms for the functions $f(r)$ and $g(r)$ aside from the asymptotic flatness condition, which requires $f(r)\,,g(r)\to 1$ as $r\to\infty$.
The resulting rotating geometries can be written as the form [13]
@@
In this work, we regard the metric (2) as the starting one to analyze the Einstein equations and get the orthonormal frame. Therefore, we do not assume the predefined static limit (3) but try to get the solution from the general physical requirements. We investigate the structure of these solutions and discuss their implications for black hole physics.
@@ 2 Rotating geometries from the static spherically symmetric black hole
Then, we require the condition that the stress tensor does not have the off-diagonal [^3] Note however that this non-vanishing value does not imply the existence of shear viscosity between the two directions. One could find the orthonormal basis which makes this value vanish by rotating $r$ and $\theta$ directions. $r\theta$ component, $T_{r\theta}=0$. This is required for fluids undergoing only a rotational motion about a fixed axis (the $z$-axis here), which leads to $R_{r\theta}=0$ [73]. To calculate the Einstein tensor, we take the orthonormal non-coordinate basis from the metric (6):
@@
$ds^{2}=-\frac{\Delta^{s}(r)}{r^{2}}dt^{2}+\frac{r^{2}}{\Delta^{s}(r)}dr^{2}+r^{2}d\Omega_{(2)}^{2}\,$, (13)
where $\Delta^{s}=r^{2}-2Mr+v(r)$. The source developing this metric satisfies $T_{\mu\nu}=G_{\mu\nu}/8\pi$ with the Einstein tensor given in Eq. (12). Therefore, we interpret that the metric (10) describes the rotating geometry of this static one.
@@
    where erf(x)(x) denotes the error function. This is an interesting generalization of the Kerr family with a minor modification of the equation of state. Note that $v(r)-rv^{\prime}(r)=v_{2}e^{-v_{c}r^{2}/2}$. Therefore, the stress tensor exponentially decreases with $r$. We analyze the properties of the solution in the next section.
@@
Starting from the metric ansatz (10), we impose the axially rotating condition $R_{\hat{r}\hat{\theta}}=0$ and select an appropriate equation of state to determine the final form of the metric.
@@ 3 Analysis of the new solution
We now determine the locations of the ergosphere and the event horizon of the rotating black hole. The ergosphere is the region located between the static limit surface and the outer event horizon [78]. The static limit surface is determined at the location $g_{tt}=0$, while the event horizon corresponds to a Killing horizon [79], which is located where $\Delta(r)|_{r_{H}}=0$. Due to the presence of a transcendental error function, precise determination of the horizon’s position becomes challenging. Therefore, we will numerically investigate it by assigning values to the coefficients $v_{2}$ and $v_{c}$, as well as the rotation parameter $a$.
Before proceeding with the numerical analysis of the event horizon, it is necessary to examine the constraints on the coefficient $v_{2}$ required for the event horizon to exist. The function $\bar{\Delta}(y)$ monotonically decreases from one at $y=0$ to zero as $y\to\infty$. Expanding $\Delta(r)$ around $r\sim 0$ and $r\gg v_{c},M$, we get
@@
Let us set the radius $r_{c}$ such that $\Delta^{\prime}(r_{c})=0$ . At this point, we have
@@
Figure 1: (color online). The characteristic form of the function $\Delta(r)$ for various values of $v_{2}$ and $a$, respectively.
Figure 1 illustrates the shape of $\Delta(r)$ as the function of the coordinate $r$. Figure 1(a) shows the shape of $\Delta(r)$ while varying the value of $v_{2}$, and Fig. 1(b) shows the shape while varying the value of the rotation parameter $a$. We take $M=1$ for simplicity. In Fig. 1(a), the gray curve represents a Schwarzschild black hole with $a=0$ and $v_{2}=0$, while the black curve represents a Kerr black hole with $v_{2}=0$. The red, orange, and green curves represent $\Delta(r)$ with $v_{2}=0.5$, $v_{2}=1.0$, and $v_{2}=1.5$, respectively, with $a=0.3$. The size of the event horizon is largest for a Schwarzschild black hole. The second largest case is the outer event horizon of a rotating Kerr black hole. When the rotation parameter $a$ held fixed, increasing $v_{2}$ leads to a reduction in the horizon size, bringing the black hole closer to the extremal limit. In Fig. 1(b), we take $v_{2}=0.5$. The blue dashed line represents $\bar{\Delta}(\sqrt{v_{c}}r)$ with $v_{c}=0.2$. The black curve represents $\Delta(r)$ with $a=0.0$, corresponding to the black hole with anisotropic matter. The red, orange, and green curves represent $\Delta(r)$ with $a=0.3$, $a=0.5$, and $a=0.6$, respectively. When the rotation parameter $a$ is set to zero, the outer horizon reaches its maximum size. As $a$ increases, the horizon size shrinks and approaches the extremal limit.
@@
Note that the presence of the $v_{2}\bar{\Delta}(y)$ term always increases the value of $\Delta(r)$. Therefore, the radius of the outer horizon $r_{+}$ (Henceforth, $r_{H}$ will be used to denote $r_{+}$.), given by $\Delta(r)|_{r_{H}}=0$, decreases because of this term. This implies that the $\bar{\Delta}(y)$ term makes the entropy of the black hole decreases, and as a result, the Kerr black hole maximizes the entropy.
Given the asymptotic values the mass $M$ and the angular momentum parameter $a$, the horizon area varies with $v_{c}$. As a result of this, the thermodynamic law can no longer be described by the asymptotic values.
In the $a\to 0$ limit, the metric reduces to Eq. (13) with $\Delta^{s}(r)=r^{2}-2Mr+v_{2}\bar{\Delta}(y)$, as given in Eq. (19), this corresponds to a new static black hole geometry with $w_{r}=-1$, and $w_{\theta}=w_{\phi}=1+v_{c}r^{2}/2$. In other words,
@@
Figure 2: (color online). The energy density as a function of $r$.
Figure 2 illustrates the profile of $\rho^{s}(r)$ as a function of $r$, with parameters $M=1$ and $v_{c}=0.2$. For the Schwarzschild black hole case, the horizon is located at $r_{H}=2$. For black holes accompanying matter fields, however, the presence of the error function complicates the precise determination of the event horizon. To address this, we identify the horizon as the point where $\Delta^{s}(r)$ vanishes, numerically approximated to within $10^{-5}$. The red, orange, and green curves represent $\rho^{s}(r)$ with $v_{2}=0.2$, $v_{2}=0.3$, $v_{2}=0.4$ respectively. As $v_{2}$ representing the contribution of matter field increases, the position of the outer horizon clearly shifts inward.
Now, let us turn to the mass energy. We begin with the case of a charged black hole. The mass energy contained within a sphere of radius $R>r_{H}$ is $m_{c}(R)=M-\frac{Q^{2}}{2R}$ [80]. The mass energy enclosed within the horizon, as viewed from the outside, is related to the irreducible mass $M_{I}$: $m_{c}(r_{H})=M-\frac{Q^{2}}{2r_{H}}=\frac{r_{H}}{2}\equiv M_{I}$, which implies $M=M_{I}+\frac{Q^{2}}{2r_{H}}$. The second term corresponds to the contribution from the charge $Q$  [81]. In summary, the energy contained within the horizon is equivalent to the portion of a black hole’s total mass energy that cannot be extracted from the black hole. For the present static black hole, The mass energy within the horizon is
$m(r_{H})=M-\frac{v(r_{H})}{2r_{H}}\,$, (21)
where $v(r_{H})=v_{2}\left[e^{-\frac{r^{2}_{H}v_{c}}{2}}+\sqrt{\frac{\pi v_{c}}{2}}r_{H}\left(\mbox{erf}\left(\sqrt{\frac{v_{c}r^{2}_{H}}{2}}\right)-1\right)\right]$. This expression has the same structure as in the charged black hole case, namely $m(r_{H})=\frac{r_{H}}{2}\equiv M_{I}$, so that $M=M_{I}+\frac{v(r_{H})}{2r_{H}}$. For the present rotating black hole, the mass energy relation from the horizon information of the black hole is given by
$M^{2}=\left[M_{Ir}+\frac{v(r_{H})}{4M_{Ir}}\right]^{2}+\frac{J^{2}}{4M^{2}_{Ir}}\,$, (22)
where $J=Ma$ denotes the angular momentum and $M_{Ir}\equiv\frac{\sqrt{r^{2}_{H}+a^{2}}}{2}$ denotes the irreducible mass of the rotating black hole. For more information on other types of anisotropic matter fields, see [20].
@@ 4 Thermodynamics
We now analyze the shape of the temperature curve for $T_{H}\geq 0$. At large $r_{H}\gg(a^{2}+v_{2}e^{-v_{c}r^{2}_{H}/2})$, the temperature behaves as $T_{H}\propto r^{-1}_{H}$, similar to the Schwarzschild black hole. Conversely, for small $r_{H}$, the term $(a^{2}+v_{2}e^{-v_{c}r^{2}_{H}/2})$ in the numerator dominates. In this regime, the temperature initially decreases as $T_{H}\propto-(r^{-3}_{H})$ transitioning to $T_{H}\propto-(r^{-1}_{H})$ as $r_{H}$ grows slightly larger. As $r_{H}\to 0$, $T_{H}$ diverges to negative infinity. There exists a specific value of $r_{H}$ where $T_{H}=0$, corresponding to an extremal black hole. Additionally, the $r_{H}$ value at which the black hole temperature reaches its maximum is determined by the condition $dT_{H}/dr_{H}=0$.
@@
$M$ $=\frac{1}{2}(2T_{H}S+2\Omega_{H}J+\Phi_{1}C_{1}+\Phi_{2}C_{2})$ (25)
$+\frac{1}{2}\sqrt{(2T_{H}S+2\Omega_{H}J+\Phi_{1}C_{1}+\Phi_{2}C_{2})^{2}-4\Omega_{H}J(2\Phi_{1}C_{1}+\Phi_{2}C_{2})}\,$,
where $\Omega_{H}=\frac{a}{r^{2}_{H}+a^{2}}$, $J=Ma$, $S=\pi(r^{2}_{H}+a^{2})$. When $v_{2}\to 0$, it reduces to the case of the Kerr black hole with $M_{K}=2T_{H}S+2\Omega_{H}J$.
@@ 5 Summary and discussions
We used the results of the Newman-Janis (NJ) algorithm [10] and the “non-complexification procedure” [13] as a method of separation of variables in the metric to solve the Einstein equation and obtain the reduced form of the metric (2). Then, we presented a new solution to the Einstein equations describing a rotating black hole as an extension of the Kerr spacetime. We first demand the Einstein tensor component $R_{r\theta}$ vanishes so that the geometry is consistent with the axial symmetry. We then, check the Einstein tensor to get proper equation of states the stress tensor should follow.
However, obtaining a rotating black hole solution becomes more challenging when the NJ algorithm is not well-defined for a given geometry. This issue arises, for instance, when $f(r)\neq g(r)$, when a cosmological constant is present, or when the black hole is immersed in a magnetic field. In the case of $f(r)\neq g(r)$, the standard approach is to solve Einstein’s equations directly by using a metric ansatz, and verify whether $R_{r\theta}=0$. An example of this method applied within the NJ algorithm framework can be found in [87]. For black holes with a cosmological constant, one may follow Carter’s method [88], while black holes immersed in the magnetic field require analysis based on Ernst spacetime [89, 90].